OpenStudy (anonymous):

What is the implicit differentiation of xsiny+cos2y=cosy

OpenStudy (anonymous):

Try and solve it yourselves, taking help from the step below. 1. differentiate the whole equation with respect to x, taking y as constant. name it 'U' 2. repeat step one but this time differentiate with respect to y taking x as constant. name it as 'V' 3. Dy/Dx = - (U/V)

myininaya (myininaya):

You need to know a few rules: (y)'=y' (x)'=1 (fg)'=f'g+fg' (cos(x))'=-sin(x) (sin(x))'=cos(x) (f(g(x)))'=g'(x) * f '(g(x))

OpenStudy (anonymous):

i tried it and i got -siny + xcos(y) dx/dy - sin (2y) * 2 dy/dx but im so confused on what to do next

OpenStudy (anonymous):

Follow my steps.. Assume Y is a constant.. then differentiate for x.. then assume X as constant.. read my reply dude..!

OpenStudy (anonymous):

ah i got it. thanks!